# Prime and Reach: Synthesising Body Motion for Gaze-Primed Object Reach

### v1

@@ Abstract
Human motion generation is a challenging task that aims to create realistic motion imitating natural human behaviour. We focus on the well-studied behaviour of priming an object/location for pick up or put down – that is, the spotting of an object/location from a distance, known as gaze priming, followed by the motion of approaching and reaching the target location. To that end, we curate, for the first time, 23.7K gaze-primed human motion sequences for reaching target object locations from five publicly available datasets, i.e., HD-EPIC, MoGaze, HOT3D, ADT, and GIMO.
We pre-train a text-conditioned diffusion-based motion generation model, then fine-tune it conditioned on goal pose or location, on our curated sequences. Importantly, we evaluate the ability of the generated motion to imitate natural human movement through several metrics, including the ‘Reach Success’ and a newly introduced ‘Prime Success’ metric. On the largest dataset, HD-EPIC (LABEL:fig:egoallo_examples), our model achieves 60% prime success and 89% reach success when conditioned on the goal object location.
@@ 1 Introduction
Text-guided human motion generation has seen significant advancements, generating increasingly realistic body motions [tevet2023mdm, guo2023momask, Guo2022t2m]. These works excelled at translating textual descriptions into a wide array of human movements, ranging from simple actions like walking and running to more complex motions such as dancing. The scope has then expanded to include navigating virtual environments towards a spatial target location [diomataris2024wandr, karunratanakul2023gmd, xie2024omnicontrol]. More recently, models are designed to generate human motions of interactions with various objects [taheri2022goal, paschalidis20253d, tendulkar2023flex, xu2023interdiff, diller2024cg], such as grasping small objects or interacting with furniture. However, these works heavily rely on synthetic datasets [tendulkar2023flex, araujo2023circle] or datasets collected in lab settings [taheri2020grab, jiang2023full, lv2024himo]. As a result, they do not capture or model the prime and reach (P&R) behaviour, synthesising motions that are less natural and thus less useful as a digital replica of human behaviour.
@@
On the other hand, egocentric data collection provides a compelling alternative solution, enabling the capture of more natural interactions directly from the first-person perspective, accompanied by a gaze-mounted camera. Several datasets capture natural interactions in diverse daily activities [ma2024nymeria, grauman2024ego] and human-human interactions [zhang2022egobody], as observed from an egocentric perspective. Nevertheless, most egocentric datasets [Damen2022RESCALING, perrett2025HD-EPIC, grauman2022ego4d] do not capture full-body motion, as pairing egocentric and 3D sensors is costly and challenging. Several recent works [li2023ego, yi2025egoallo, guzov-jiang2025hmd2, chi2025estimating] have explored estimating full-body motion conditioned on the egocentric camera pose and viewpoint, training on diverse data [ma2024nymeria, grauman2024ego]. Although all these egocentric datasets capture diverse human behaviour, they have not been explored for modelling the preparatory priming and reaching motion (see Fig LABEL:fig:egoallo_examples for sample sequences).
@@
Human visuomotor coordination is fundamentally anticipatory, continuously integrating sensory information to facilitate fluid, goal-directed actions [jia2025learning]. The anticipatory mechanism heavily relies on gaze. By directing visual attention to relevant objects or areas before reaching, gaze offers crucial predictive signals that enable the motor system to prepare and execute actions efficiently and naturally [land1999roles, johansson2001eye, hayhoe2005eye, jia2025learning]. The ability to prime and reach has recently been investigated in various real-world applications, including robotics [kerrj2025eyerobot, rahrakhshan2022learning, shafti2019gaze]. For example, in a manipulation involving a towel, the robot would visually fixate on the towel and the grasp point before extending its arm [kerrj2025eyerobot]. However, such behaviour has not been explored in training or evaluating motion generation.
@@
In this work, we enable the synthesis (or generation) of priming and reaching behaviour by combining egocentric datasets that offer gaze priming and full-body motion with conditioned diffusion models.
We list our contributions as follows:
  • We explore, for the first time, synthesising full-body motion to replicate humans’ natural ability to first gaze-prime then approach and reach objects for interactions.
  • We curate five datasets, totalling 23.7K prime and reach sequences from five public datasets.
  • We train goal-conditioned motion generation models on our curated datasets and evaluate their ability to generate prime and reach motion. We compare conditioning on the goal pose vs the goal object location.
  • We introduce the ‘Prime Success’ metric to particularly evaluate the ability to synthesise gaze priming behaviour.
@@
  • Our results show that using goal pose as a condition, our model can boost priming ability by up to 1818x compared to previous methods, and can similarly improve reach success by up to 88x. When conditioning on the target object location only, our model improves the priming ability by up to 55x.
@@ 2.1 Human Motion Generation
Human motion generation aims to create realistic, continuous human movements that simulate or animate natural human motion. The majority of work in this field generates human motion from a single modality such as text [Guo2022t2m, petrovich22temos, petrovich23tmr, zhong2023attt2m, athanasiou2022teach], action [guo2020action2motion, cervantes2022implicitneuralrepresentationsvariable, petrovich21actor, posegpt], speech [zhu2023taming, alexanderson2023listen], and music [li2021learn, tseng2022edge, sun2022nonfreezing], with a few recent works tackle motion generation from multiple modalities [li2024unimotion, genmo2025, bian2024motioncraft, luo2024m3gpt].
Text-to-motion Generation.
@@
Initial efforts [ahuja2019language2pose, ghosh2021synthesis] in text-to-motion were deterministic, converging to an averaged motion given an input text. After the advent of the denoising diffusion models [ho2020denoising, song2021denoising], these models are nowadays a common practice to generate text-conditioned human motion [tevet2023mdm, zhang2024motiondiffuse, zhang2023remodiffuse, kim2023flame]. MDM [tevet2023mdm], MotionDiffuse [zhang2024motiondiffuse], and FLAME [kim2023flame] are conditioned on features extracted by a pre-trained text encoder. Another common approach is to disentangle motion representation from generation [guo2023momask, jiang2024motiongpt, chuan2022tm2t, zhang2023t2mgpt]. This is a two-stage process: first, a VQ-VAE [van2017neural] is trained to create a discrete codebook that tokenises continuous motion sequences. Then, a Transformer-based autoregressive model is trained on these discrete tokens to learn the motion primitives by predicting the next token in a sequence. Although the text can serve as a strong signal for conditioning semantic motion, these methods often lack precise control over body positions.
Location-Conditioned Motion Generation.
@@
Our work is related to a line of research that generates human motion conditioned on a target location. Several works extend the text-to-motion model to enable various spatial controls (e.g., entire root trajectory [karunratanakul2023gmd], keyframe root location [karunratanakul2023gmd], obstacle avoidance [karunratanakul2023gmd], temporally and spatially sparse joints [pinyoanuntapong2025maskcontrol, xie2024omnicontrol], or goal location [tevet2025closd]). Recently, WANDR [diomataris2024wandr] introduced a data-driven model conditioned on the initial pose and the goal location of the right wrist to generate avatars that walk and reach the goal in 3D space. Other works [li2024egogen, diomataris2025movinglookingvisiondrivenavatar] address the goal-reaching human motion generation via egocentric perception and reinforcement learning. Despite these advancements, these methods rely on synthetic datasets [araujo2023circle] or MoCap-based datasets [mahmood2019amass], which limit their ability to generate natural interactions in real-world scenarios. Additionally, these works do not address or evaluate priming. In this work, we curate the first set of datasets that include full-body, priming, and reaching, with a focus on replicating this human priming-then-reaching motion through generation.
Ego-body Pose Generation.
@@
Recent research has explored estimating [jiang2022avatarposer, jiang2024egoposer, Luo2021DynamicsRegulatedKP, chi2025estimating, guzov-jiang2025hmd2] or forecasting [yuan2019ego, escobar2025egocast, patel25uniegomotion, Hatano2025EgoH4] human motion from an egocentric perspective. These methods typically adopt a generative approach as the human body is largely invisible from an egocentric view, unlike ego-body pose estimation from a downward-facing camera [Millerdurai_EventEgo3D_2024, wang2024egocentric, hakada2022unrealego, tome2019xr, cuevas2024simpleego, zhao2021egoglass]. EgoEgo [li2023ego] is the first work to propose head pose (camera pose) conditioned human motion generation, but was mainly evaluated on synthetic datasets. Subsequently, EgoAllo [yi2025egoallo] proposes a head-centric representation (i.e., canonicalisation) to achieve spatial and temporal invariance, and also enables the integration of in-view 3D hand poses for better prediction. We utilise the ego-body pose estimation method [yi2025egoallo] to generate human motion on gaze-primed and reach sequences curated from egocentric datasets.
2.2 Eye-gaze in Motion
@@
Eye-gaze is an important predictive signal that directs attention and primes the processing of future movements [land1999roles, johansson2001eye, hayhoe2005eye]. Recognising the critical role of gaze, recent research has been focusing on estimating gaze/saliency [lai2022eye, lai2024eye, chong2018connecting] or explicitly leveraging this cue for various problems, such as video understanding [mazzamuto2025gazing, peng2025eyemllmbenchmarkingegocentric] or human-robot interactions [stolzenwald2018can, saran2018human]. Several works focus on future motion prediction following gaze priming [wei2018and, yan2023gazemodiff, hu2025hoigaze, lou2024multimodal, jia2025learning]. Tian et al. [tian2024gaze] generate hand-object interactions but only in table-top settings. Different from these works, we wish to synthesise both the gaze priming and the reach motion, for the full body, conditioned on the goal.
@@ 3 Prime and Reach Data Curation
We first introduce the principle of curating ‘Prime and Reach’ (P&R) sequences from longer videos. We then detail the steps we carried out to curate these sequences from five public datasets. We note the statistics of these sequences, which we use for training and evaluation.
@@ 3.1 P&R sequence Curation
Interaction datasets include multiple and frequent object reaching and manipulation behaviours. However, a critical aspect largely unexplored is the role of gaze in priming or “spotting” objects prior to the reaching motion. We take this missed opportunity and curate for the first time P&R motion sequences from datasets capturing wearable gaze and object interactions. We are inspired by the “gaze priming” discussed in [perrett2025HD-EPIC] where objects were annotated in 3D and then used to identify the fixation that occurs prior to the physical action, signalling intent of interaction.
@@
Specifically, we project the user’s gaze into the 3D environment to form a ray. First, the gaze direction provided by the eye-tracker in the camera’s local coordinate system at any time $t$, $\mathbf{p}^{t}_{\text{gaze\_cam}}$, is transformed using the camera-to-world transformation matrix $\mathbf{T}^{t}_{\text{c2w}}$. The final normalised gaze direction vector, $\hat{\mathbf{d}}^{t}_{\text{gaze}}$, is then computed as the vector from the camera’s world position, $\mathbf{o}^{t}_{\text{cam}}$, to this new world-space gaze point $\mathbf{p}^{t}_{\text{gaze\_world}}$ as shown in Equation 1.
@@
We register a relevant location as primed if the gaze ray, originating from the camera’s position, intersects with the corresponding 3D bounding box or 3D location $o_{\text{3D}}$. Therefore, we define the prime time $t_{p}$ as
@@
where $T_{\text{int}}$ is the set of all timestamps within the temporal window $t_{e}-w,t_{e}]$, where the gaze ray intersects the 3D location and $t_{p}$ is the first moment where the intersection happens. We discard sequences where $T_{\text{int}}=\emptyset$. Following [perrett2025HD-EPIC], we use $w=10$ secs. To compute the intersection i.e. $\text{intersect}(\hat{\mathbf{d}}^{t}_{\text{gaze}},o_{\text{3D}})$, we use the slab test method [Majercik2018Voxel], details of which are available in the supplementary. At the end of this process, we get P&R sequences each defined by a prime time $t_{p}$ and reach time $t_{e}$.
Table 1: Curated Dataset Statistics. We report statistics on curated P&R sequences across five publicly available datasets, ordering them by the size of curated sequences. We report the number of P&R sequences, duration between prime time and reach time i.e. $t_{e}-t_{r}$ (Prime Gap), body pose type, the distance/movement of body and hand. * indicates that body poses are estimated using [yi2025egoallo].
@@
HD-EPIC [perrett2025HD-EPIC] 18,134 $5.49\pm 2.76$ $3.55\pm 2.79$ SMPL-H* $0.72\pm 0.67$ $0.45\pm 0.22$
MoGaze [kratzer2020mogaze] 2,637 $3.64\pm 0.94$ $1.53\pm 0.92$ 3D Skeleton $1.07\pm 0.62$ $0.75\pm 0.25$
HOT3D [banerjee2025hot3d] 2,416 $4.31\pm 1.54$ $2.37\pm 1.57$ SMPL-H* $0.20\pm 0.15$ $0.38\pm 0.19$
ADT [pan2023aria] 411 $7.44\pm 2.47$ $4.51\pm 2.74$ 3D Skeleton $1.23\pm 1.28$ $0.56\pm 0.24$
GIMO [zheng2022gimo] 130 $7.11\pm 2.49$ $4.47\pm 1.49$ SMPL-X [pavlakos2019expressive] $3.09\pm 1.20$ $0.64\pm 0.19$
@@ 3.2 Datasets
As explained in Sec. 3.1, we formulate how P&R sequences can be curated from long video sequences. We consider five publicly available human-object interaction datasets, all of which contain 2D gaze tracked from wearable gaze trackers along with camera poses [perrett2025HD-EPIC, kratzer2020mogaze, banerjee2025hot3d, pan2023aria, zheng2022gimo].
@@
MoGaze [kratzer2020mogaze] is human motion data designed explicitly for human-object interactions, with a particular focus on ‘pick’ and ‘put’. The dataset includes synchronised full-body motion captured using motion capture markers, 3D object models, and 6-DoF and eye-gaze data. The dataset contains 180 minutes of motion capture data from seven participants performing pick-and-put actions along with temporal segment annotations of these actions. Using the gaze data and the object locations, we determine priming timestamps ($t_{p}$) for each pick and put event. We slice the motion data for $t\in[t_{p}-2,t_{e}]$ constructing $2,637$ P&R sequences.
@@
HOT3D [banerjee2025hot3d] captures 3D hand-object interactions. The dataset offers 198 ARIA recordings featuring 14 subjects interacting with 33 diverse objects. We only use the ARIA videos as these provide gaze information. As pick/put timestamp annotations are not provided, we extract temporal segments where an object is in-hand by thresholding (<5<5 cm) the distance between the nearest hand vertices and the object locations. After identifying these in-hand segments, we refine the segment boundaries by detecting the object state change from stationary to in-hand or vice-versa. This gives us pick/put events $t_{e}$. We use gaze and camera pose to estimate the prime time for these events, resulting in $2,416$ P&R sequences. We estimate full-body motion for these sequences using EgoAllo [yi2025egoallo].
@@
Aria Digital Twin (ADT) [pan2023aria] provides a rich collection of synchronised data, including images, eye-tracking data, 6-DoF object data, and 3D human poses. 72 videos in the dataset capture indoor activities and interactions involving 398 unique objects and provide paired eye gaze and 3D body motion data. We curate P&R sequences from these videos. Same as HOT3D, we find temporal segments when objects are in-hand by thresholding the distance between object locations and nearest wrist locations from the corresponding body poses. We identify pick and put events near the temporal boundaries of these segments based on how the object state changed. These events were then primed to determine $t_{p}$, resulting in 411 P&R sequences with full-body motion data.
@@
GIMO [zheng2022gimo] is a benchmark that focuses on intent-guided human motion segments. It provides 217 trimmed segments, along with corresponding SMPL-X fitted IMU-captured body poses and egocentric views with eye gaze data captured by the HoloLens 2. We discard all segments that include resting activities e.g., sitting or lying on the bed. Gimo segments include object pick-up but no put-down. However, they do not provide the 3D locations or timestamps of object pick-up. Therefore, we manually annotate these event timestamps ($t_{e}$) from RGB videos and use the relevant wrist location at the timestamp as our object locations. We determine $t_{p}$ for each of these events following the same method as before. We get 130 P&R SMPL-X body motion sequences.
@@
In total, we curate $23,728$ P&R sequences from the five datasets. Statistics are provided in Tab. 1. We showcase sample P&R sequences in Fig. 2. Importantly, we unify body pose formats across datasets by representing them using the canonicalised 22-joint body motion used in HumanML3D [Guo_2022_CVPR]. Following [tevet2023mdm], we convert the 22 joint positions to 263-dim vector representation that combines local pose, rotation and velocity of each joint. The curated sequences from each dataset are split 70%-30% into train and test sets. Details are provided in supplementary.

### v2

# : Equal Contribution https://masashi-hatano.github.io/prime-and-reach/

@@ Abstract
Human motion generation is a challenging task that aims to create realistic motion imitating natural human behaviour. We focus on the well-studied behaviour of priming an object/location for pick up or put down – that is, the spotting of an object/location from a distance, known as gaze priming, followed by the motion of approaching and reaching the target location. To that end, we curate, for the first time, 23.7K gaze-primed human motion sequences for reaching target object locations from five publicly available datasets, i.e., HD-EPIC, MoGaze, HOT3D, ADT, and GIMO. We pre-train a text-conditioned diffusion-based motion generation model, then fine-tune it conditioned on goal pose or location, on our curated sequences. Importantly, we evaluate the ability of the generated motion to imitate natural human movement through several metrics, including the ‘Reach Success’ and a newly introduced ‘Prime Success’ metric. Tested on 5 datasets, our model generates diverse full-body motion, exhibiting both priming and reaching behaviour, and outperforming baselines and recent methods.
Figure 1: Prime & Reach sequences from HD-EPIC [67], using full-body pose from EgoAllo [94]. (Left) A sequence starting with the intention to reach the container (cyan sphere). Gaze priming is evident -gaze (green plus) intersecting the object- during the approach before reaching the object. (Right) Similar behaviour is noted for priming and picking up the scale (cyan sphere). [darker colors indicate later time].
@@ 1 Introduction
Text-guided human motion generation works excel at translating textual descriptions into a wide array of human movements [84, 22, 23], from simple actions like running to complex motions such as dancing. The scope has expanded to include navigating environments [15, 39, 91] and interactions with objects [80, 63, 82, 92, 13]. However, these works heavily rely on synthetic datasets [82, 4] or controlled in lab recordings [81, 36, 56]. Consequently, they fail to model natural behaviours like prime and reach (P&R), limiting their utility as digital human replicas.
@@
On the other hand, egocentric data collection provides a compelling alternative solution, enabling the capture of diverse daily activities [57, 21] and human-human interactions [99] accompanied by a head-mounted eye-tracking camera. Nevertheless, most egocentric datasets [12, 67, 20] do not capture full-body motion, as pairing egocentric and 3D sensors is costly and challenging. Several recent works [48, 94, 27, 9] have explored estimating full-body motion conditioned on the egocentric camera pose and viewpoint, training on diverse data [57, 21]. Egocentric datasets have not been explored for generative modelling, including the preparatory priming and reaching motion (see Fig 1 for sample sequences).
@@
Human visuomotor coordination is fundamentally anticipatory, continuously integrating sensory information to facilitate fluid, goal-directed actions [32]. The anticipatory mechanism heavily relies on gaze. By directing visual attention to relevant objects or areas before reaching, gaze offers crucial predictive signals that enable the motor system to prepare and execute actions efficiently and naturally [45, 37, 29, 32]. The ability to prime and reach has recently been investigated in various real-world applications, including robotics [40, 72, 76]. For example, in a manipulation involving a towel, the robot would visually fixate on the towel and the grasp point before extending its arm [40]. However, such behaviour has not been explored in training or evaluating motion generation.
In this work:
@@
  • We enable, for the first time, the synthesis (or generation) of priming and reaching behaviour by combining egocentric datasets that offer gaze priming and full-body motion with conditioned diffusion models.
  • We curate 23.7K prime and reach sequences from five datasets, by leveraging gaze and object-location annotations to automatically extract motion segments that contain priming and reaching.
  • We design goal-conditioned models generating prime and reach motions, introducing a ‘Prime Success’ metric to evaluate against existing baselines.
@@
  • Using goal pose as a condition, our model can boost priming ability by up to $18.2\%$ absolute gain over the previous best-performing method, while achieving reach success nearly perfectly. When conditioning on the object location, our model improves the priming ability by up to $19.5\%$.
@@ 2 Related Work
Human motion generation aims to create realistic, continuous human movements that simulate or animate natural human motion. The majority of work in this field generates human motion from a single modality such as text [23, 69, 70, 103, 5], action [26, 8, 68, 53], speech [104, 3], and music [50, 87, 79], with a few recent works tackling motion generation from multiple modalities [46, 49, 7, 54].
@@
Text-to-motion Generation. Initial efforts [1, 19] in text-to-motion were deterministic, converging to an averaged motion given an input text. After the advent of the denoising diffusion models [30, 77], these models are nowadays a common practice to generate text-conditioned human motion [84, 97, 98, 41]. MDM [84], MotionDiffuse [97], and FLAME [41] are conditioned on features extracted by a pre-trained text encoder. Another common approach is to disentangle motion representation from generation [22, 33, 25, 96]. This is a two-stage process: first, a VQ-VAE [88] is trained to create a discrete codebook that tokenises continuous motion sequences. Then, a Transformer-based autoregressive model is trained on these discrete tokens to learn motion primitives by next-token-prediction. Although the text can serve as a strong signal for conditioning semantic motion, these methods often lack precise control over body positions.
@@
Location-Conditioned Motion Generation. Our work is related to a line of research that generates controllable human motion [71, 91, 38, 83, 39, 101, 15]. Guided Motion Diffusion (GMD) [39] extends text-to-motion models with spatial controls, including root trajectories, keyframe locations, obstacle avoidance, and sparse joint constraints. DNO [38] instead treats a pre-trained text-to-motion diffusion model as a motion prior and optimises the initial noise vector with task-specific gradients (e.g., goal joints or target location) at inference time, enabling flexible control without retraining. Furthermore, DartControl [101] proposed autoregressive prediction with latent motion diffusion conditioned on history sequences and textual description. It achieves motion-in-between via DNO and goal-reaching task through policy control. Recently, WANDR [15] introduced a data-driven model conditioned on the initial pose and the goal location of the right wrist to generate avatars that walk and reach the goal in 3D space. Other works [47, 14] address the goal-reaching human motion generation via egocentric perception and reinforcement learning. Despite these advancements, these methods rely on synthetic datasets [4] or MoCap-based datasets [58], which limit their ability to generate natural interactions in real-world scenarios. Additionally, these works do not address or evaluate priming. In this work, we curate the first set of datasets that include full-body, priming, and reaching, with a focus on replicating this human priming-then-reaching motion through generation.
@@
Ego-body Pose Generation. Recent research has explored estimating [35, 34, 55, 9, 27] or forecasting [95, 18, 64, 28] human motion from an egocentric perspective. These methods typically adopt a generative approach as the human body is largely invisible from an egocentric view, unlike ego-body pose estimation from a downward-facing camera [61, 89, 2, 86, 11, 100]. EgoEgo [48] is the first work to propose head pose (camera pose) conditioned human motion generation, but was mainly evaluated on synthetic datasets. Subsequently, EgoAllo [94] proposes a head-centric representation (i.e., canonicalisation) to achieve spatial and temporal invariance, and also enables the integration of in-view 3D hand poses for better prediction. We utilise this method to generate human motion on gaze-primed and reach sequences curated from egocentric datasets.
@@
Eye-gaze in Motion. Eye-gaze is an important predictive signal that directs attention and primes the processing of future movements [45, 37, 29]. Recognising the critical role of gaze, recent research has been focusing on estimating gaze/saliency [43, 44, 10] or explicitly leveraging this cue for various problems, such as video understanding [60, 66] or human-robot interactions [78, 75]. Several works focus on future motion prediction following gaze priming [90, 93, 31, 52, 32]. Tian et al. [85] generate hand-object interactions but only in table-top settings. Different from these works, we wish to synthesise both the gaze priming and the reach motion, for the full body, conditioned on the goal.
@@ 3 Prime and Reach Data Curation
We first introduce the principle of curating ‘Prime and Reach’ (P&R) sequences from longer videos (Sec.˜3.1). We then detail the steps we carried out to curate these sequences from five public datasets (Sec.˜3.2). We note the statistics of these sequences, which we use for training and evaluation.
@@ 3.1 P&R sequence Curation
Interaction datasets include multiple and frequent object reaching and manipulation behaviours. However, a critical aspect largely unexplored is the role of gaze in priming or “spotting” objects prior to the reaching motion. We take this missed opportunity and curate for the first time P&R motion sequences from datasets capturing wearable gaze and object interactions. We are inspired by the “gaze priming” discussed in [67] where objects were annotated in 3D and then used to identify the fixation that occurs prior to the physical action, signalling intent of interaction.
@@
[…]
@@
[…]
@@
where $T_{\text{int}}$ is the set of all timestamps within the temporal window $t_{e}-w,t_{e}]$, where the gaze ray intersects the 3D location and $t_{p}$ is the first moment where the intersection happens. We discard sequences where $T_{\text{int}}=\emptyset$. Following [67], we use $w=10$ secs. To compute the intersection i.e. $\text{intersect}(\hat{\mathbf{d}}^{t}_{\text{gaze}},o_{\text{3D}})$, we use the slab test method [59], details of which are available in the supplementary. At the end of this process, we get P&R sequences each defined by a prime time $t_{p}$ and reach time $t_{e}$.
Table 1: Curated Dataset Statistics. We report statistics on curated P&R sequences across five publicly available datasets, ordering them by the size of curated sequences. We report the number of P&R sequences, duration between prime time and reach time i.e. $t_{e}-t_{p}$ (Prime Gap), body pose type, the distance/movement of body and hand. * indicates that body poses are estimated using [94].
@@
HD-EPIC [67] 18,134 $5.49\pm 2.76$ $3.55\pm 2.79$ SMPL-H* [51] $0.72\pm 0.67$ $0.45\pm 0.22$
MoGaze [42] 2,637 $3.64\pm 0.94$ $1.53\pm 0.92$ 3D Skeleton $1.07\pm 0.62$ $0.75\pm 0.25$
HOT3D [6] 2,416 $4.31\pm 1.54$ $2.37\pm 1.57$ SMPL-H* [51] $0.20\pm 0.15$ $0.38\pm 0.19$
ADT [62] 411 $7.44\pm 2.47$ $4.51\pm 2.74$ 3D Skeleton $1.23\pm 1.28$ $0.56\pm 0.24$
GIMO [102] 130 $7.11\pm 2.49$ $4.47\pm 1.49$ SMPL-X [65] $3.09\pm 1.20$ $0.64\pm 0.19$
@@ 3.2 Datasets
As explained in Sec.˜3.1, we formulate how P&R sequences can be curated from long video sequences. We consider five publicly available human-object interaction datasets, all of which contain 2D gaze tracked from wearable gaze trackers along with camera poses [67, 42, 6, 62, 102].
@@
MoGaze [42] is human motion data designed explicitly for human-object interactions, with a particular focus on ‘pick’ and ‘put’. The dataset includes synchronised full-body motion captured using motion capture markers, 3D object models and their 6-DoF, and eye-gaze data. The dataset contains 180 minutes of motion capture data from seven participants performing pick-and-put actions along with temporal segment annotations of these actions. Using gaze and object locations, we determine priming timestamps ($t_{p}$) for each event. We slice the motion data for $t\in[t_{p}-2,t_{e}]$ constructing $2,637$ P&R sequences.
@@
HOT3D [6] captures 3D hand-object interactions. The dataset offers 198 Aria recordings featuring 14 subjects interacting with 33 diverse objects. We only use the Aria videos as these provide gaze information. As pick/put timestamp annotations are not provided, we extract temporal segments where an object is in-hand by thresholding (<5<5 cm) the distance between the nearest hand vertices and the object locations. After identifying these in-hand segments, we refine the segment boundaries by detecting the object state change from stationary to in-hand or vice-versa. This gives us pick/put events $t_{e}$. We use gaze and camera pose to estimate the prime time for these events, resulting in $2,416$ P&R sequences. We estimate full-body motion for these sequences using EgoAllo [94].
@@
Aria Digital Twin (ADT) [62] provides a rich collection of synchronised data, including images, eye-tracking data, 6-DoF object data, and 3D human poses. 72 videos in the dataset capture indoor activities and interactions involving 398 unique objects and provide paired eye gaze and 3D body motion data. We curate P&R sequences from these videos. Same as HOT3D, we find temporal segments when objects are in-hand by thresholding the distance between object locations and nearest wrist locations from the corresponding body poses. We identify pick and put events near the temporal boundaries of these segments based on how the object state changed. These events were then primed to determine $t_{p}$, resulting in 411 P&R sequences with full-body motion data.
@@
GIMO [102] is a benchmark that focuses on intent-guided human motion segments. It provides 217 trimmed segments, along with corresponding SMPL-X fitted IMU-captured body poses and egocentric views with eye gaze data captured by the HoloLens 2. We discard all segments that include resting activities without pick/put actions e.g., sitting or lying on the bed. GIMO segments include object pick-up but no put-down. However, they do not provide the 3D locations or timestamps of object pick-up. Therefore, we manually annotate these event timestamps ($t_{e}$) from RGB videos and use the relevant wrist location at the timestamp as our object locations. We determine $t_{p}$ following the same method and get 130 P&R SMPL-X body motion sequences.
@@
In total, we curate $23,728$ P&R sequences from the five datasets. Statistics are provided in Tab.˜1 and sample P&R sequences in Fig.˜2. Importantly, we unify body pose formats across datasets by representing them using the canonicalised 22-joint body motion used in HumanML3D [24]. Following [84], we convert the 22 joint positions to 263-dim vector representation that combines local pose, rotation and velocity of each joint. The curated sequences are split 70%-30% into train and test sets.